# The Polar Express: Optimal Matrix Sign Methods and Their Application to the Muon Algorithm

### v1

@@ 4.4   Finite precision considerations
When working in finite-precision arithmetic, especially the half-precision bfloat16 format used in deep learning, we must take some care to avoid blowups and other problems due to numerical error. To this end, we make three small changes to the method. These adjustments stabilize the algorithm with a negligible effect on accuracy. Observe that these adjustments can be made in the offline stage by modifying the coefficients of our optimal polynomials.
The first issue arises when numerical round-off creates singular values that are slightly larger than $u_{t}$. Our optimal polynomials converge only when the singular values of $\bm{X}_{t}$ are actually less than $u_{t}$. In some cases we have $p_{t}(u_{t}+\epsilon)>u_{t+1}+\epsilon$, so over many iterations, a singular value that is slightly too large could end up growing to $\infty$ instead of converging to $1$. (This issue is unlikely to arise for degree $3,7,11,15\ldots$, for which $p_{t}(u_{t})=\ell_{t+1}<u_{t+1}$). To fix this issue, we simply replace each polynomial $x\mapsto p_{t}(x)$ by $x\mapsto p_{t}(x/1.01)$. This safety factor corrects for round-off errors in previous iterations while only slightly changing the behavior of the polynomial on the interval $[\ell_{t},u_{t}]$, though it does cause the singular values to converge to $0.999998$ instead of to $1$. To correct for this, the safety factor can be omitted in the final iteration.
The second issue was identified in [35] and addressed in the context of polynomial iterations by Chen and Chow [12]. In general, iterative methods for $\operatorname*{polar}(\bm{M})$ aim to increase each singular value relative to the largest singular value; while $\sigma_{\min}(\bm{X}_{0})\ll\sigma_{\max}(\bm{X}_{0})$, after enough iterations, $\sigma_{\min}(\bm{X}_{t})\approx\sigma_{\max}(\bm{X}_{t})\approx 1$. However, the convergence of each singular value to $\sigma_{\max}$ may not be monotonic. Over the domain $\ell_{t}\ll u_{t}$, our optimal polynomial $p_{t}$ oscillates repeatedly between $\ell_{t+1}$ and $u_{t+1}$, so some singular values that are near $u_{t}$ may get mapped down to $\ell_{t+1}$. It so happens that this non-monotonicity—even of a single singular value at a single iteration—causes loss of precision. That is, problems occur if
@@
where $0\leq\sigma_{\min}\leq\sigma_{i}\leq\sigma_{\max}$ are singular values of $\bm{X}_{t}$ [35]. (In the extreme case, if $p_{t}(\sigma_{i})<0$, the $i$th singular vector will change sign and the method will converge to the polar factor of the wrong matrix.) Unlike Newton-Schulz, unscaled Newton, or QDWH, our method is affected by this loss of precision. To mitigate this issue, [12] modify their update polynomials as follows to impose a lower limit on $\frac{p_{t}(\sigma_{i})}{\sigma_{i}}$. Notice that the issue only occurs when $\ell_{t}\ll u_{t}$; as $\ell_{t}\to u_{t}$, our optimal polynomial approaches the Padé approximant and so $\frac{p_{t}(x)}{x}\geq 1$ for all $x\in[0,u_{t}]$. We could fully solve the problem by using the Padé approximant instead of our optimal polynomial, but this would significantly slow convergence. Instead we compromise. When $\ell_{t}\geq u_{t}/10$, we find that $\frac{p_{t}(x)}{x}\geq 0.231$. Therefore, whenever $\ell_{t}<u_{t}/10$ we select the update rule as though $\ell_{t}=u_{t}/10$. This change slows convergence, but only very slightly. (The choice of 10 is somewhat arbitrary. In Appendix G, we use a different factor.)
The third change is copied from the original Muon implementation: normalize $\bm{M}$ by $\|\bm{M}\|_{\text{F}}+10^{-7}$ instead of by $\|\bm{M}\|_{\text{F}}$. As before, we set $u_{1}=1$.
@@ 4.5   The algorithm
We give the complete pseudocode for our proposed method in Algorithm 1. Our algorithm first computes the polynomials $p_{1},\ldots,p_{T}$ of Theorem 4.1 in full precision using the Remez algorithm. This stage is offline because the coefficients of the polynomials are only computed and stored once. For every subsequent call to the algorithm, these coefficients are reused and the offline stage is skipped. The polynomial $p^{\star}:=p_{T}\circ\cdots\circ p_{1}$ is then applied to the input matrix $\bm{M}$ in the online stage. The online stage can be performed in lower precision (bfloat16) for greater speed on a GPU.
Algorithm 1 The Polar Express
@@
Horner’s rule can be used to carry out each iteration. For instance, if $p_{t}=ax+bx^{3}+cx^{5}$, then
@@
where $\bm{Y}_{t-1}=\bm{X}_{t-1}^{\top}\bm{X}_{t-1}$. A simple implementation of the offline stage of Algorithm 1 is given in Appendix G.
For deep learning applications, we recommend using $d=5$ and $T=5$ or $6$ with $\ell_{1}=10^{-3}$. With these parameters, the offline stage gives the following sequence of polynomials, as generated by the code in Appendix G:
$\begin{split}p_{1}(x)&=8.20516x-22.90193x^{3}+16.46072x^{5}\\ p_{2}(x)&=4.06692x-2.86128x^{3}+0.51838x^{5}\\ p_{3}(x)&=3.91349x-2.82425x^{3}+0.52485x^{5}\\ p_{4}(x)&=3.30601x-2.43023x^{3}+0.48695x^{5}\\ p_{5}(x)&=2.30402x-1.64272x^{3}+0.40091x^{5}.\end{split}$ (17)
All told, our proposal for Muon is to apply the composition of the these polynomials to $\bm{M}/(\|\bm{M}\|_{F}+10^{-7})$.
@@ 5.1   Convergence of Polar Express
We compare the performance of Polar Express against degree-5 Newton-Schulz, and the methods of Chen and Chow, Jordan, and You. We first study an idealized scenario where the spectrum of the input matrix is known exactly. We generate a random matrix whose singular values are evenly spaced on a logarithmic scale between $10^{-6}$ and $1$. The right and left singular vectors are chosen at random. The left panel of Figure 4 shows the results. Since all the methods in this plot use degree-5 polynomials, their computational and runtime costs are all proportional to the number of iterations. As expected, Newton-Schulz converges but makes almost no progress for the first 17 iterations. Jordan’s method achieves error of $\approx 0.3$ after just 11 iterations, but ceases to converge further. You’s method, which is only defined for six iterations, is barely visible on this plot. When Polar Express is instantiated with $\ell=\sigma_{\min}$, it dominates the other methods at every iteration, achieving excellent accuracy after just 11 iterations and converging about twice as fast as Newton-Schulz to any given error. Even when the lower bound on $\sigma_{\min}$ is wrong by two orders of magnitude in either direction, the method remains competitive, though it does not actually outperform Keller until iteration 13 or 14.
@@
We next test the methods’ performance on a matrix from a real-world application, namely, the gradient of a weight matrix from the fourth transformer block of a GPT-2 architecture with respect to a language modeling objective on a batch of text from the Tiny Shakespeare dataset [25]. The right panel of Figure 4 shows the results. Once again, the best-tuned version of Polar Express outperforms the other methods. This time, we see that setting $\ell$ to be many orders of magnitude too small can delay convergence significantly.
For most other weight matrices in this GPT-2 model, the methods all take more than 10 iterations to converge in the spectral norm. The spectral error is be large if there is even one outlying singular value that is far from $1$. However, for some applications, we may be satisfied with a weaker notion of convergence, like the relative Frobenius norm. Figure 5 shows the performance of various methods on this metric. We use gradient matrices of the same model, but from two different layers. In addition, we compare the degree-5 methods to Chen and Chow’s degree-3 method. To make this comparison fair, we measure the number of matrix-matrix products performed by each method instead the number of iterations. We find that Polar Express can once again dominate the other methods across iterations. Chen and Chow’s method is also quite competitive, and the remaining methods behave much as in Figure 4.
@@ 5.2   Training GPT-2
In our final experiment, we compare the performance of using our Polar Express method as given in (17) inside the Muon algorithm versus Jordan’s [23] and You’s [10] methods.
Our experimental setup is based on the modified nanogpt code of Jordan [22]. We train a GPT-2 model with $n_{\text{embd}}=768$, $n_{\text{layer}}=12$, $n_{\text{head}}=12$, and a vocabulary size of $50{,}257$, using a context length of $1024$. Training is performed on 1B tokens from the FineWeb dataset [3], using a batch size of 32 and a single epoch. All models are trained with mixed precision (bfloat16) on 4 H100 GPUs. For all methods we use the learning rate schedule proposed in [22], consisting of a constant phase for the first 40% of training steps followed by a linear decay. All methods for the matrix sign computations are performed in float16b precision and use five iterations.
@@
Figure 6 shows the resulting runs of each method in terms of validation loss and training loss, where we can see that muon-PolarExp, achieves better validation and training loss than muon-Jordan or muon-You. Since each iteration of the different matrix sign methods are equally expensive (since they all apply a degree 5 polynomial), improved validation loss in terms of epochs also translates to an improved loss in terms of wall clock time (see bottom right of Figure 6). The advantage is remarkably consistent across all learning rates and epochs.
@@
The aim of this section is to prove Theorem 4.1. We begin with a result that provides a few essential properties for the the polynomial solving (8) when $T=1$. This result is known as Chebyshev’s theorem [11] or the equioscillation theorem [45, Chapter 10].
@@
Then Algorithm 1 returns $\bm{X}_{T}=p^{\star}(\bm{M})$. Let $h\in\mathbb{P}_{q}$ be $[q/0]$ Padé-approximant to $(1-x)^{-1/2}$ [26, Section 3] and define $p(x)=xh(1-x^{2})\in\mathbb{P}_{d}^{\operatorname*{odd}}$. Define $f=p\circ\cdots\circ p$ as the composition of $p$ with itself $T$ times. Then, by Theorem 4.1 and [26, Theorem 3.1] we have
@@ Appendix C Proof of equivalence between (7) and (8)
$=\max_{x\in[\ell,u]}\left|1-p(x)\right|.$
Consequently, $\varepsilon_{1}\geq\varepsilon$. Hence, $\varepsilon_{1}=\varepsilon_{2}$, as desired.
@@ Appendix D Remez algorithm
Let $d=2q+1$. Recall that from Lemma A.1 the optimal polynomial must satisfy the equioscillation property $q+2$ times. The Remez algorithm is an iterative algorithm that finds the equioscillation points $A=\{x_{0},\ldots,x_{q+1}\}$ from Lemma A.1 by iteratively refining a sequence of trial points $A^{(k)}=\{x_{0}^{(k)},\ldots,x_{q+1}^{(k)}\}$ so that $A^{(k)}$ converges to $A$. From the sequence of trial points $A^{(k)}$ the algorithm also finds a sequence of polynomials $p^{(k)}$ so that $p^{(k)}$ converges to the optimal polynomial. The convergence is extremely fast, and usually 10 iterations is sufficient to converge to the optimal polynomial up to double precision machine epsilon [38]. More commonly, the Remez algorithm is used to find optimal polynomial approximations to general continuous functions where $d\approx 100$ or even $d\approx 1000$. However, because the polynomial we build to approximate $\operatorname*{sign}(x)$ is a composition of polynomials, each of which has a low degree, in our setting the degree $d$ is small, usually $d=5$. For $d=5$ the Remez algorithm admits a simple description, as we outline below. The following description can be generalized to arbitrary $d$.
Recall that as stated in Lemma A.1, the unique optimal approximation $p^{\star}\in\mathbb{P}_{5}^{\operatorname*{odd}}$ satisfies the equioscillation property four times. The Remez algorithm first starts with a trial set $A^{(1)}=\{x_{0}^{(1)},x_{1}^{(1)},x_{2}^{(1)},x_{3}^{(1)}\}\subset[\ell,u]$ which ideally should approximately satisfy the equioscillation property. Since we know that $\ell$ and $u$ must be equioscillation points we always set $x_{0}^{(k)}=\ell$ and $x_{3}^{(k)}=u$ for all $k$. $x_{2}^{(1)}$ and $x_{3}^{(1)}$ are chosen equispaced to be $\frac{1}{4}\ell+\frac{3}{4}u$ and $\frac{3}{4}\ell+\frac{1}{4}u$, since we observe that as $\ell\approx u$ these are approximately the equioscillation points. Next, the algorithm solves the following system of equations
@@
Once we have solved for $a_{1},b_{1},c_{1},$ and $E_{1}$ we set $p^{(k)}(x)=a_{1}x+b_{1}x^{3}+b_{1}x^{5}$. Now we want to find the worst case error of using $p^{(k)}$ to approximate $x\mapsto 1$. We therefore find local maxima of the error function $e_{1}(x)=1-p_{1}(x)$ on $(\ell,u)$ by setting the derivative of $e_{1}(x)$ to zero and solving for $x$. This results in solving a quadratic equation $05b_{1}x^{4}+3b_{1}x^{2}+a=$ 0, which have closed form solutions from the quadratic formula. We now let $x_{1}^{(2)}$ and $x_{2}^{(2)}$ be the solutions to this equation and let $A^{(2)}=\{\ell,x_{1}^{(2)},x_{2}^{(2)},u\}$ and repeat the procedure until $|E_{k}|:=\max\limits_{x\in[\ell,u]}|1-p^{(k)}(x)|\approx\max\limits_{x\in[\ell,u]}|1-p^{(k+1)}(x)|:=|E_{k+1}|$.
We emphasize that the matrix appearing in (20) is a Vandermonde matrix, which become notoriously ill-conditioned as $d$ grows large [16, Section 4.6]. However, since in our setting we allow $d$ to be small there is no ill-conditioning due to the size of the matrix. Instead, we observe ill-conditioning when $\ell\approx u$. However, as $\ell/u\to 1$ the optimal polynomial will converge to the polynomial $\frac{x/u}{8}\left(15-10(x/u)^{2}+3(x/u)^{4}\right)$, which can be verified by noting that as $\ell/u\to 1$ all equioscillation points $x_{0},x_{1},x_{2},x_{3}$ must converge to $u$. For general $d=2q+1$, the polynomial will converge to $(x/\ell)h(1-(x/\ell)^{2})$ where $h\in\mathbb{P}_{q}$ is the $[q/0]$ Padé approximant to $(1-x)^{1/2}$ [26]. In fact, this polynomial is extremely close to the optimal polynomial for sufficiently large $\ell$. To see this, let $p^{\star}$ be the optimal approximation from $\mathbb{P}_{5}^{\operatorname*{odd}}$ and let $p(x)=\frac{x/u}{8}\left(15-10(x/u)^{2}+3(x/u)^{4}\right)$. Then,
@@
where we invoked [26, Theorem 3.1] and the fact that $p^{\star}$ is the optimal approximation to $x\mapsto 1$ from $\mathbb{P}_{5}^{\operatorname*{odd}}$. Hence, when $\ell/u\geq 1-\epsilon_{d}^{1/3}$, where $\epsilon_{\text{double}}\approx 1.1\times 10^{-16}$ is the double precision machine epsilon, then $|p^{\star}(x)-p(x)|\leq 2\epsilon_{\text{double}}$. In other words, up to double precision machine epsilon, $p^{\star}$ is equal to $p$. Therefore, whenever $\ell/u\geq 1-\epsilon_{\text{double}}^{1/3}$ the algorithm simply returns $p$ as the optimal polynomial.

### v2

@@ 4.4   Finite precision considerations
When working in finite-precision arithmetic, especially the half-precision bfloat16 format used in deep learning, we must take some care to avoid blowups and other problems due to numerical error. To this end, we make three small changes to the method. These adjustments stabilize the algorithm with a negligible effect on accuracy. Furthermore, these adjustments can be made in the offline stage by modifying the coefficients of our optimal polynomials.
The first issue arises when numerical round-off creates singular values that are slightly larger than our current upper bound $u_{t}$. Our optimal polynomials converge only when the singular values of $\bm{X}_{t}$ are less than $u_{t}$. In some cases we have
$p_{t}(u_{t}+\epsilon)>u_{t+1}+\epsilon,$
so over many iterations, a singular value that is slightly larger than $u_{t}$ large could grow to $\infty$ instead of converging to $1$.
To fix this issue, we simply replace each polynomial $x\mapsto p_{t}(x)$ by $x\mapsto p_{t}(x/1.01)$. This safety factor corrects for round-off errors in previous iterations while only slightly changing the behavior of the polynomial on the interval $[\ell_{t},u_{t}]$, though it does cause the singular values to converge to $0.999998$ instead of to $1$. To correct for this, the safety factor can be omitted in the final iteration.
The second issue was identified in [34] and addressed in the context of polynomial iterations by Chen and Chow [11]. In general, iterative methods for $\operatorname*{polar}(\bm{M})$ aim to increase each singular value relative to the largest singular value; while $\sigma_{\min}(\bm{X}_{0})\ll\sigma_{\max}(\bm{X}_{0})$, after enough iterations, $\sigma_{\min}(\bm{X}_{t})\approx\sigma_{\max}(\bm{X}_{t})\approx 1$. However, the convergence of each singular value to $\sigma_{\max}$ may not be monotonic. Over the domain $[\ell_{t},u_{t}]$, our optimal polynomial $p_{t}$ oscillates repeatedly between $\ell_{t+1}$ and $u_{t+1}$, so some singular values that are near $u_{t}$ may get mapped down to $\ell_{t+1}$. It so happens that this non-monotonicity—even at a single iteration—can cause loss of precision. That is, problems occur if
@@
where $0\leq\sigma_{\min}\leq\sigma_{i}\leq\sigma_{\max}$ are singular values of $\bm{X}_{t}$ [34]. In the extreme case $p_{t}(\sigma_{i})<0$, the $i$th singular vector will change sign, casuing the method to converge to the polar factor of the wrong matrix. Unlike Newton-Schulz, unscaled Newton, or QDWH, our method is affected by this loss of precision.
To mitigate this issue, [11] propose modifying their update polynomials to enforce a lower bound on the ratio $\frac{p_{t}(\sigma_{i})}{\sigma_{i}}$. This issue only occurs when $\ell_{t}\ll u_{t}$; as $\ell_{t}\to u_{t}$, our optimal polynomial approaches the Padé approximant and so $\frac{p_{t}(x)}{x}\geq 1$ for all $x\in[0,u_{t}]$. We could fully solve the problem by using the Padé approximant instead of our optimal polynomial, but this would significantly slow down convergence. Instead we compromise. When $\ell_{t}\geq u_{t}/10$, we find that $\frac{p_{t}(x)}{x}\geq 0.236$. Therefore, whenever $\ell_{t}<u_{t}/10$ we select the update rule as though $\ell_{t}=u_{t}/10$. This change slows convergence, but only very slightly. (The choice of 10 is somewhat arbitrary. In Appendix G, we use a different factor.)
The third change is copied from the original Muon implementation: normalize $\bm{M}$ by $\|\bm{M}\|_{\text{F}}+10^{-2}$ instead of by $\|\bm{M}\|_{\text{F}}$. As before, we set $u_{1}=1$.
@@ 4.5   The algorithm
We give the complete pseudocode for our proposed method for any degree in Algorithm 2. We give the specfic version of the Polar Express with degree $d=5$ and $\ell=10^{-3}$ used in our GPT experiments in Algorithm 1. Our algorithm first computes the polynomials $p_{1},\ldots,p_{T}$ of Theorem 4.1 in full precision using the results of Section 4.2 (or the Remez algorithm for $d>5$). This stage is offline because the coefficients of the polynomials are only computed and stored once. For every subsequent call to the algorithm, these coefficients are reused and the offline stage is skipped. For instance, in Algorithm 1 these polynomials have been precomputed and stored in the variable coeffs_list.
Algorithm 2 The General Polar Express
@@
The polynomial $p^{\star}:=p_{T}\circ\cdots\circ p_{1}$ is then applied to the input matrix $\bm{M}$ in the online stage. The online stage can be performed in lower precision (bfloat16) for greater speed on a GPU. Horner’s rule can be used to carry out each iteration. For instance, if $p_{t}=ax+bx^{3}+cx^{5}$, then
@@
where $\bm{Y}_{t-1}=\bm{X}_{t-1}^{\top}\bm{X}_{t-1}$.
A simple implementation of the offline stage of Algorithm 2 is given in Appendix G. For deep learning applications, we recommend using $d=5$ and $T=5$ or $6$ with $\ell_{1}=10^{-3}$. With these parameters, the offline stage as implemented in Appendix G gives the polynomials encoded in coeffs_list in Algorithm 1. All told, our proposal for Muon is to apply the composition of these polynomials to $\bm{M}/(\|\bm{M}\|_{F}+10^{-2})$.
@@ 5.1   Convergence of Polar Express
We compare the performance of Polar Express against degree-5 Newton-Schulz and the methods of Chen and Chow [11], Jordan [22], and You [9].
We first study an idealized scenario where the spectrum of the input matrix is known exactly. We generate a random matrix whose singular values are evenly spaced on a logarithmic scale between $10^{-6}$ and $1$. The right and left singular vectors are chosen at random. The left panel of Figure 4 shows the results. Since all the methods in this plot use degree-5 polynomials, their computational and runtime costs are all proportional to the number of iterations. As expected, Newton-Schulz converges but makes almost no progress for the first 17 iterations. Jordan’s method rapidly achieves an error of $\approx 0.3$ after just 11 iterations, but ceases to converge further. You’s method, which is difficult to see on the plot because it is only defined for six iterations, converges at a similar rate as Jordan’s method. When Polar Express is instantiated with $\ell=\sigma_{\min}$, it dominates the other methods at every iteration, achieving excellent accuracy after just 11 iterations and converging about twice as fast as Newton-Schulz to any given error. Even when the lower bound on $\sigma_{\min}$ is wrong by two orders of magnitude in either direction, the method remains competitive, though it does not actually outperform Keller until iteration 13 or 14.
@@
Next we test the methods’ performance on a matrix from a real-world application, namely, the gradient of a weight matrix from the fourth transformer block of a GPT-2 architecture with respect to a language modeling objective on a batch of text from the Tiny Shakespeare dataset [24]. The right panel of Figure 4 shows the results. Once again, the best-tuned version of Polar Express outperforms the other methods. This time, we see that setting $\ell$ to be many orders of magnitude too small can delay convergence significantly, and make Polar Express less competitive as compared to Jordan’s method.
For most other weight matrices in this GPT-2 model, the methods all take more than 10 iterations to converge in the spectral norm. The spectral error is large if there is even one outlying singular value that is far from $1$. However, for some applications, we may be satisfied with a weaker notion of convergence, like the relative Frobenius norm. Figure 5 shows the performance of various methods on this metric. We use gradient matrices of the same model, but from two different layers. In addition, we compare the degree-5 methods to Chen and Chow’s degree-3 method. To make this comparison fair, we measure the number of matrix-matrix products performed by each method instead the number of iterations. We find that Polar Express can once again dominate the other methods across iterations. Chen and Chow’s method is also quite competitive, and the remaining methods behave much as in Figure 4.
@@ 5.2   Training GPT-2
In our final experiment, we compare the performance of using our Polar Express method given in Algorithm 1 inside the Muon algorithm versus Jordan’s [22] and You’s [9] methods.
Our experimental setup is based on the modified nanogpt code of Jordan [21]. We train two different GPT-2 models:
[…]
and a vocabulary size of $50{,}257$, using a context length of $1024$. Training is performed on 1B tokens from the FineWeb dataset [2], using a batch size of 32 and a single epoch. All models are trained with mixed precision (bfloat16) on 4 H100 GPUs. For all methods we use the learning rate schedule proposed in [21], consisting of a constant phase for the first 40% of training steps followed by a linear decay. All methods for the matrix sign computations are performed in float16b precision and use five iterations.
We apply Muon selectively to certain layers of the model. Following the nano-gpt implementation [21], we assign Muon to all parameters with at least two dimensions (typically weight matrices, and excluding RMS norm parameters), excluding the embeddings, unembeddings, and the attention head. These excluded parameters are instead optimized with AdamW.
@@
Figure 1 and Figure 6 shows the resulting runs of each method in terms of validation loss and training loss on the GPT-Large and GPT-Small models, respectively. In both figures we can see that muon-PolarExp achieves a better validation and training loss than muon-Jordan or muon-You for every learning rate. Since each iteration of the different matrix sign methods are equally expensive (since they all apply a degree 5 polynomial), improved validation loss in terms of epochs also translates to an improved loss in terms of wall clock time (see bottom right of Figure 6). The advantage is remarkably consistent across all learning rates and epochs.
@@
The aim of this section is to prove Theorem 4.1. We begin with a result that provides a few essential properties for the the polynomial solving (7) when $T=1$. This result is known as Chebyshev’s theorem [10] or the equioscillation theorem [46, Chapter 10].
@@
Then Algorithm 2 returns $\bm{X}_{T}=p^{\star}(\bm{M})$. Let $h\in\mathbb{P}_{q}$ be $[q/0]$ Padé-approximant to $(1-x)^{-1/2}$ [25, Section 3] and define $p(x)=xh(1-x^{2})\in\mathbb{P}_{d}^{\operatorname*{odd}}$. Define $f=p\circ\cdots\circ p$ as the composition of $p$ with itself $T$ times. Then, by Theorem 4.1 and [25, Theorem 3.1] we have
@@ Appendix C Proof of equivalence between (6) and (7)
$=\max_{x\in[\ell,u]}\left|1-p(x)\right|\;=\varepsilon_{2}$
Consequently, $\varepsilon_{1}\geq\varepsilon_{2}$. Hence, $\varepsilon_{1}=\varepsilon_{2}$, as desired.
@@ Appendix D Remez algorithm
Let $d=2q+1$. Recall that from Lemma A.1 the optimal polynomial must satisfy the equioscillation property $q+2$ times. The Remez algorithm is an iterative algorithm that finds the equioscillation points $A=\{x_{0},\ldots,x_{q+1}\}$ from Lemma A.1 by iteratively refining a sequence of trial points $A^{(k)}=\{x_{0}^{(k)},\ldots,x_{q+1}^{(k)}\}$ so that $A^{(k)}$ converges to $A$. From the sequence of trial points $A^{(k)}$ the algorithm also finds a sequence of polynomials $p^{(k)}$ so that $p^{(k)}$ converges to the optimal polynomial. The convergence is extremely fast, and usually 10 iterations is sufficient to converge to the optimal polynomial up to double precision machine epsilon [37]. More commonly, the Remez algorithm is used to find optimal polynomial approximations to general continuous functions where $d\approx 100$ or even $d\approx 1000$. However, because the polynomial we build to approximate $\operatorname*{sign}(x)$ is a composition of polynomials, each of which has a low degree, in our setting the degree $d$ is small, usually $d=5$. For $d=5$ the Remez algorithm simplifies significantly. We now describe this simplified algorithm.
Recall that as stated in Lemma A.1, the unique optimal approximation $p^{\star}\in\mathbb{P}_{5}^{\operatorname*{odd}}$ satisfies the equioscillation property at four points. The Remez algorithm first starts with a trial set $A^{(1)}=\{x_{0}^{(1)},x_{1}^{(1)},x_{2}^{(1)},x_{3}^{(1)}\}\subset[\ell,u]$ which ideally should come close to satisfying the equioscillation property. Since we know that $\ell$ and $u$ must be equioscillation points we always set $x_{0}^{(k)}=\ell$ and $x_{3}^{(k)}=u$ for all $k$. $x_{2}^{(1)}$ and $x_{3}^{(1)}$ are chosen to be $\frac{1}{4}\ell+\frac{3}{4}u$ and $\frac{3}{4}\ell+\frac{1}{4}u$, since we observe that as $\ell\to u$ these are approximately the other equioscillation points. Next, the algorithm solves the following system of equations
@@
Once we have solved for $a_{1},b_{1},c_{1},$ and $E_{1}$ we set $p_{1}(x)=a_{1}x+b_{1}x^{3}+b_{1}x^{5}$. Now we want to find the error in the $L^{\infty}$ norm of using $p_{1}$ to approximate $x\mapsto 1$. We therefore find local maxima of the error function $e_{1}(x)=1-p_{1}(x)$ on $(\ell,u)$ by setting the derivative of $e_{1}(x)$ to zero and solving for $x$. This results in the quartic equation $05b_{1}x^{4}+3b_{1}x^{2}+a=$ 0, which has closed form solutions given by the quadratic formula after the substitution $y=x^{2}$. We now let $x_{1}^{(2)}$ and $x_{2}^{(2)}$ be the solutions to this equation and let $A^{(2)}=\{\ell,x_{1}^{(2)},x_{2}^{(2)},u\}$. We repeat the procedure until $|E_{k}|:=\max\limits_{x\in[\ell,u]}|1-p_{k}(x)|\approx\max\limits_{x\in[\ell,u]}|1-p_{k+1}(x)|:=|E_{k+1}|$.
We note that the matrix appearing in (18) is a Vandermonde matrix. Vandermonde matrices become notoriously ill-conditioned as the degree grows large [15, Section 4.6]. However, since in our setting we choose $d$ to be small, there is no ill-conditioning. Instead, we observe ill-conditioning when $\ell\approx u$. However, as $\ell/u\to 1$ the optimal polynomial will converge to the polynomial $\frac{x/u}{8}\left(15-10(x/u)^{2}+3(x/u)^{4}\right)$, which can be verified by noting that as $\ell/u\to 1$ all equioscillation points $x_{0},x_{1},x_{2},x_{3}$ must converge to $u$. For general $d=2q+1$, the polynomial will converge to $(x/\ell)h(1-(x/\ell)^{2})$ where $h\in\mathbb{P}_{q}$ is the $[q/0]$ Padé approximant to $(1-x)^{1/2}$ [25]. In fact, this polynomial is extremely close to the optimal polynomial for sufficiently large $\ell$. To see this, let $p^{\star}$ be the optimal approximation from $\mathbb{P}_{5}^{\operatorname*{odd}}$ and let $p(x)=\frac{x/u}{8}\left(15-10(x/u)^{2}+3(x/u)^{4}\right)$. Then,
@@
where we invoked [25, Theorem 3.1] and the fact that $p^{\star}$ is the optimal approximation to $x\mapsto 1$ from $\mathbb{P}_{5}^{\operatorname*{odd}}$. Hence, when $\ell/u\geq 1-\epsilon_{d}^{1/3}$, where $\epsilon_{\text{double}}\approx 1.1\times 10^{-16}$ is the double precision machine epsilon, then $|p^{\star}(x)-p(x)|\leq 2\epsilon_{\text{double}}$. In other words, up to double precision machine epsilon, $p^{\star}$ is equal to $p$. Therefore, whenever $\ell/u\geq 1-\epsilon_{\text{double}}^{1/3}$ the algorithm simply returns $p$ as the optimal polynomial.